OpenStudy (anonymous):

Using Calculus, fi nd two positive numbers with sum of 100 and product as large as possible. (Answer: both numbers are 50, what are the steps?)

OpenStudy (anonymous):

using x and y as the variables a good idea? x + y = 100?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

looking for max(x*y), and an equation to represent that?

OpenStudy (welshfella):

now make y the subject

OpenStudy (anonymous):

y = 100 - x

OpenStudy (welshfella):

and form an eqauation for the area in terms of x

OpenStudy (anonymous):

area = x * (100 - x)

OpenStudy (anonymous):

You have that x + y = 100, and you want to maximize K = x*y, then we have K = (100-x)*x differentiating and making the equal to 0: 100-2x = 0 ==> x = 50, now find y

OpenStudy (welshfella):

right now find the derivative da/dx

OpenStudy (anonymous):

DA = 1*(100-x) + x(0 - 1)

OpenStudy (welshfella):

A = 100x - x^2 dA/dx = 100 - 2x = 0 2x = 100 x = 50 and finally y = 100-50 == 50

OpenStudy (anonymous):

so, where y = 50 is the global maximum of that graph, which means the maximum area possible

OpenStudy (anonymous):

thanks for the help